# ConRebSeg: A Segmentation Dataset for Reinforced Concrete Construction

### v1

@@ Abstract
The construction industry has been traditionally slow in adopting digital technologies. However, these are becoming increasingly necessary due to a plentitude of challenges, such as a shortage of skilled labor and decreasing productivity levels compared to other industries. Autonomous robotic systems can alleviate this problem, but the software development process for these systems is heavily driven by data, a resource usually challenging to find in the construction domain due to the lack of public availability. In our work, we therefore provide a dataset of 14,805 RGB images with segmentation labels for reinforced concrete construction and make it publicly available. We conduct a detailed analysis of our dataset and discuss how to deal with labeling inconsistencies. Furthermore, we establish baselines for the YOLOv8L-seg, DeepLabV3, and U-Net segmentation models and investigate the influence of data availability and label inconsistencies on the performance of these models. Our study showed that the models are precise in their predictions but would benefit from more data to increase the number of recalled instances. Label inconsistencies had a negligible effect on model performance, and we, therefore, advocate for a crowd-sourced dataset to boost the development of autonomous robotic systems in the construction industry.
Keywords dataset  $\cdot$ construction robotics  $\cdot$ segmentation  $\cdot$ rebar detection  $\cdot$ shotcrete  $\cdot$ digitization
@@ 1 Introduction
The construction industry will be facing a big challenge in the upcoming years. In the European Union (EU), the number of people employed in the construction industry has steadily increased since 2015 [1]. However, so did the number of hours worked per person, with the exception of a little, probably COVID-related dip in 2020 [2]. Both of these statistics suggest that the construction industry will be facing major labor shortages in the future. In addition, the real labor productivity per hour worked has decreased compared to 2015 [2]. This particular trend does not apply to all industries, as the industrial sector without construction experienced an increase in real labor productivity [2]. A major difference between other industries and construction is the lack of adoption of digitization and automation in the construction industry. Construction companies see big potential in the adoption of these technologies, yet seem to have little or no implementation of these in place, according to a 2016 study conducted by Roland Berger [3]. Another study focused on the United States construction industry, conducted by McKinsey&Company, confirms this. In the study, construction is shown to be one of the sectors with the lowest adoption rate of digital technologies [4]. Furthermore, it shows a correlation between the adoption rate of digital technologies and productivity growth, where construction is attested with the most negative growth rate between 2005 and 2015 [4].
@@
One of the key challenges faced during the development and deployment of autonomous robotic systems in construction is the nature of the environment the systems need to operate in. For example, manufacturing environments, where robotic systems are widely adopted nowadays, are structured environments with clear, designated spaces for people, objects, etc. In addition, a manufacturing process’s outcome is usually known in advance, making it easier to program robots to follow specific, pre-planned activities. This is generally not the case in construction and this discrepancy is shown in [7].
To counterbalance the unstructured and highly dynamic nature of construction sites, autonomous robotic systems need to rely on advanced perception capabilities, integrating sophisticated systems of sensors and algorithms to e.g. detect irregularities and navigate safely through a construction site. Most often, these detection algorithms are based on neural networks and require abundant data to train on. However, unlike other industries, the construction industry lacks this abundance of data. Our work addresses this issue and provides a new dataset focused on reinforced concrete construction processes. The dataset consists of RGB images, both from our own capturing sessions and publicly available data from YouTube. The images show different scenarios of shotcrete construction in both construction and repair phases, such as the construction of ground support walls, tunnels, and culverts, as well as the repair of piles and beams. Along with these images, we tackle the task of segmentation and provide labels for every image. The images in conjunction with the labels are then the key for training deep learning-based segmentation algorithms for solving the segmentation task, which is a crucial element in enabling robots to autonomously perform construction tasks where reinforced concrete structures are involved. Furthermore, as opposed to classic Computer Vision (CV) datasets, our dataset is sequence-based, reflecting the temporospatial nature of robotic tasks.
@@
  • An overview of recent advancements in reinforcement bar detection algorithms, construction robots, and datasets in construction.
  • A publicly available dataset of construction imagery for reinforcement concrete construction with segmentation labels for four categories: exposed reinforcement bars, persons, cars, and trucks.
  • An extensive analysis of this dataset.
  • A series of baseline experiments using deep learning-based segmentation models trained on our dataset.
@@ 2.2 Reinforcement bar detection
Rebars are an important element at all lifecycle stages of structures as they are of vital importance for their integrity. In early lifecycle stages, i.e. when structures are to be newly built, CV can be used to locate and assess the quality of rebar lattices/mats [18, 19, 20, 21], specifically the spacing and linkage between rebar elements as a crucial parameter to their contribution to the structural integrity of bridges and buildings. While lattices have a characteristic regular structure, there also exists work on locating individual bars, e.g. in precast reinforced concrete [22]. During the lifetime of a building or other infrastructure, it is important to monitor structural health. Defects can either appear on the surface in form of spallations and exposed rebars—as mentioned previously, or underneath the surface. Sub-surface structural health can be surveyed using CV techniques on Ground Penetrating Radar (GPR) data [23, 24, 25] as opposed to commonly used RGB cameras and 3D Light Detection and Ranging (LiDAR) sensors.
@@ 2.3 Construction robots
The previously mentioned related work focuses on inspection of structures and site monitoring and shows plenty of research activity in these areas. They provide a good basis for the adoption of autonomous systems, which rely on semantic understanding capabilities to navigate construction sites and carry out construction activities. Examples of such systems are excavators capable of autonomously building stone walls [26] or autonomous surface excavation [27]. Other examples are autonomous rebar tying robots [28, 29] and autonomous 3D concrete printing robots with rebar placement [30]. A pipeline detecting exposed rebars from RGB images and reconstructing a 3D model from stereo images for autonomous robotic placement of shotcrete is presented in [31]. [32] developed a proof of concept showcasing the robotic application of shotcrete, and [33] shows an algorithm to identify and reconstruct steel rebar meshes. A thorough review of the state of construction robotics can be found in [34].
@@ 2.4 Datasets
Task-specific datasets are more tailored to solving a specific construction task. Table 1 gives an overview of datasets used in the works in Section 2.2 depicting rebars for various tasks. As the table shows, availability of data is low, which poses a barrier to the development of robust, deep learning-based solutions. Not every work makes a data availability statement, and more often than not, the data is only available upon request. Only [39] provides their data without any additional barriers. We want to overcome this barrier by releasing our dataset to the general public, hoping that it will motivate other researchers to make also their data publicly available.
@@ 3.1 Overview
We have collected a dataset of 14,805 images captured mostly in construction environments. The main purpose of this dataset is to collect a broad set of images depicting exposed rebars in the context of shotcrete construction. Our dataset is sequence-based and contains sequences from various shotcrete construction tasks, such as the construction of ground support walls, tunnels, and culverts as well as the repair of piles and beams. This captures the diverse appearance of exposed rebars.
@@
  • Exposed rebars: These can be whole rebar lattices, i.e. in ground support wall construction, or partially exposed lattices and single bars in defective concrete.
  • Persons: These will usually be construction workers in PPE
  • Cars: Those are regular-sized passenger cars
  • Trucks: This is usually heavy construction machinery or material delivery trucks
In total, there are 54,115 instances, distributed as follows: 41,237 exposed rebars, 11,275 persons, 330 cars, and 1273 trucks. Our dataset is a mix of self-collected data, for which the setup is described in Section 3.2, and YouTube videos.
@@ 3.2 Acquisition of data and labeling process
The sequences in the dataset stem from two sources: self-collected data we acquired during construction site visits and data from YouTube. The camera we used in our data collection visits is a Basler a2A1920-160ucBAS RGB camera fitted with a Basler C125-0418-5M-P f4mm lens. We performed two data collection sessions for our self-captured data. A summary of the two recording sessions is provided in Table 2.
Table 2: Metadata of self-collected sequences
Location Vester Sogade, 1601 Copenhagen, DK Langebro, 1219 Copenhagen, DK
@@
Short Description Balcony repair works with exposed rebars, both individual and in lattice format Repair of bridge piles and beams in challenging lighting conditions
Number of sequences 12 14
Number of frames 4911 7270
Average frame rate 1.8 FPS 3.8 FPS
Resolution (width/height) 1920x1200 1920x1200
@@
Arranging construction site visits always relies on partners with connections to the construction industry, which can be a great barrier in the process since not every site owner is willing to let outsiders in. That is why in addition to the self-collected data, we have also identified publicly accessible videos on YouTube depicting shotcrete construction activity, which has allowed us to rapidly grow the dataset. However, the frames of those videos are not directly included in the dataset, but rather their annotations together with a mapping to a specific frame of the video. This is due to legal uncertainties regarding the distribution of downloaded YouTube videos  [^1] In their service terms, YouTube states that ”You are not allowed to: […] 1. access, reproduce, download, distribute, transmit, broadcast, display, sell, license, alter, modify or otherwise use any part of the Service or any Content except: […] (c) as permitted by applicable law;” and ”[…] access the Service using any automated means (such as robots, botnets or scrapers) except: […] (c) as permitted by applicable law;[41]. We stipulate that the EU DSM directive [42] and its national implementation in Denmark [43] are fulfilling the conditions on which users are allowed to download data from YouTube as per their service terms.. Instead, we provide a way for users to obtain the data themselves in an automated manner for scientific/research purposes. Users should check the legality of this process in their respective jurisdictions. Once we have collected the data, we annotated them to obtain masks for the previously listed object classes. To make the annotation process privacy-preserving, we have treated all sequences with the face-blurring algorithm Deface [44], based on the CenterFace [45] face detector. Furthermore, we have treated all frames of the sequences gathered at Langebro (see Table 2) with histogram equalization and AI-denoising based on NAFNet [46] to alleviate the rather low-illumination, low-contrast frames. This facilitates the annotation process and avoids low-quality annotations due to invisible image features. However, we provide the original images in our dataset to give users the freedom to pre-process the frames according to their needs.
@@
While classes such as person, truck, and car have some common sense associated with them, the ExposedBars class is rather domain-specific and not easy to describe. We set the following description for the annotation process: area with exposed rebars, where the tips of the bars should be polygon points and roughly outline the area. Along that, we gave some visual examples as illustrated in Figure 1.
@@ 3.3 Statistics
Figure 2 also shows a clear dominance of the ExposedBars class, followed by the person class. In total, there are 8570, 3543, and 2692 samples for the train, val, and test split respectively. Table 3 shows key statistics about the splits and their annotations.
@@ 3.4 Distribution and structure of the dataset
We provide a public repository that contains our data and other elements that facilitate viewing, exploring, and experimenting with the dataset. We use the FiftyOne dataset management framework [47] that provides end-users with a web interface for viewing samples and managing metadata such as train/test/val tags as well as other tags describing certain properties of samples. Our repository contains scripts that set up the dataset for the user, especially obtaining the data from YouTube as we do not provide the frames directly for reasons specified in Section 3.2. In general, the data is organized as listed in Figure 4.
@@
The angle brackets in Figure 4 are placeholders and denote the meaning of the fields. dddd are decimal seconds, ns_epoch_timestamp denotes a timestamp in nanoseconds in Unix Epoch time (since Jan 1, 1970), %04d means that frame numbers up to 1000 are zero-padded.
The YouTube frames are downloaded as whole videos and then are converted to frames, however, only every 50th frame is used since the informational value of the frames in between is minimal due to minimal variations of the depicted scene. For integrity purposes, we save an MD5 checksum for every sample in the dataset, which is saved as a metadata field in the dataset management tool. At import time, every frame is checked against this checksum to ensure the integrity of the dataset.

### v2

# Segmentation Dataset for Reinforced Concrete Construction

@@ Abstract
This paper provides a dataset of 14,805 RGB images with segmentation labels for autonomous robotic inspection of reinforced concrete defects. Baselines for the YOLOv8L-seg, DeepLabV3, and U-Net segmentation models are established. Labelling inconsistencies are addressed statistically, and their influence on model performance is analyzed. An error identification tool is employed to examine the error modes of the models. The paper demonstrates that YOLOv8L-seg performs best, achieving a validation mIOU score of up to 0.59. Label inconsistencies were found to have a negligible effect on model performance, while the inclusion of more data improved the performance. False negatives were identified as the primary failure mode. The results highlight the importance of data availability for the performance of deep learning-based models. The lack of publicly available data is identified as a significant contributor to false negatives. To address this, the paper advocates for an increased open-source approach within the construction community.
keywords:
dataset , construction robotics , segmentation , rebar detection , shotcrete , digitization
[…]
[1]organization=DTU - Technical University of Denmark, Department of Electrical and Photonics Engineering, addressline=Elektrovej Bygning 326, postcode=2800, city=Kongens Lyngby, country=Denmark
@@ 1 Introduction
The construction industry will be facing a big challenge in the upcoming years. In the European Union (EU), the number of people employed in the construction industry has steadily increased since 2015 [1]. However, so did the number of hours worked per person, except for a little, probably COVID-related dip in 2020 [2]. Both of these statistics suggest that the construction industry will be facing major labor shortages in the future. In addition, the real labor productivity per hour worked has decreased compared to 2015 [2]. This particular trend does not apply to all industries, as the industrial sector without construction experienced an increase in real labor productivity [2]. A major difference between other industries and construction is the lack of adoption of digitization and automation in the construction industry. Construction companies see big potential in the adoption of these technologies, yet seem to have little or no implementation of these in place, according to a 2016 study conducted by Roland Berger [3]. Another study focused on the United States construction industry, conducted by McKinsey&Company, confirms this. In the study, construction is shown to be one of the sectors with the lowest adoption rate of digital technologies [4]. Furthermore, it shows a correlation between the adoption rate of digital technologies and productivity growth, where construction is attested with the most negative growth rate between 2005 and 2015 [4].
@@
One of the key challenges faced during the development and deployment of autonomous robotic systems in construction is the nature of the environment the systems need to operate in. For example, manufacturing environments, where robotic systems are widely adopted nowadays, are structured environments with clear, designated spaces for people, objects, etc. In addition, a manufacturing process’s outcome is usually known in advance, making it easier to program robots to follow specific, pre-planned activities. This is generally not the case in construction, and this discrepancy is shown in [7].
To counterbalance the unstructured and highly dynamic nature of construction sites, autonomous robotic systems need to rely on advanced perception capabilities, integrating sophisticated systems of sensors and algorithms to e.g., detect irregularities and navigate safely through a construction site. Most often, these detection algorithms are based on neural networks and require abundant data to train on. However, unlike other industries, the construction industry lacks this abundance of data. Our work addresses this issue and provides a new dataset focused on reinforced concrete construction processes, specifically the inspection and repair of reinforcement bars. The dataset consists of 14,805 RGB images, both from our own capturing sessions and publicly available data from YouTube. The images show different scenarios of shotcrete construction in both construction and repair phases, such as the construction of ground support walls, tunnels, and culverts, as well as the repair of piles and beams. With these images, we tackle the task of image segmentation for robotic inspection, repair, and construction of reinforced concrete structures. The images are provided with labels, which can be used to train deep learning-based segmentation algorithms for solving the segmentation task, which is a crucial element in enabling robots to autonomously perform construction tasks where reinforced concrete structures are involved. Such autonomous systems are currently in development, for example by the RobétArmé project [8], which is developing an autonomous robotic solution to shotcrete construction and benefits from datasets like ours for training deep learning-based methods for semantic understanding of the construction environment [9]. We demonstrate that the dataset can be used to train a working segmentation model by establishing baselines for the state-of-the-art models DeepLabV3, U-Net and YOLOv8L-seg. We perform experiments analyzing the influence of data availability on the performance of YOLOv8L-seg. We show that the model can be successfully used on an embedded GPU typically found in robotic platforms. Furthermore, as opposed to classic Computer Vision (CV) datasets, our dataset is scene-based and of sequential nature, i.e., the images have a temporal order, reflecting the spatio-temporal nature of robotic tasks. This property is used to identify label inconsistencies and analyze the effect of these on the performance of the model. Based on these experiments, we identify construction-related challenges that hinder the application of these technologies in the construction domain and propose steps to alleviate them.
@@
  • 1.
    A new open source dataset of construction imagery for reinforcement concrete construction with segmentation labels for four categories: exposed reinforcement bars, persons, cars, and trucks.
  • 2.
    An analysis protocol for label inconsistencies to analyze the influence of multiple valid labelling styles on model performance, i.e., opinions of different engineers regarding the extent of a defect.
  • 3.
    Introducing emphasis on spatio-temporal properties when employing CV techniques in a construction robotics context
  • 4.
    Investigating the influence of data availability and pre-training on the performance of the YOLOv8L-seg segmentation model.
  • 5.
    A discussion about challenges faced when applying deep learning-based methods in the construction sector, based on our experimental findings.
@@ 2.2 Reinforcement bar detection
Rebars are an important element at all lifecycle stages of structures, as they are of vital importance for their integrity. In early lifecycle stages, i.e., when structures are to be newly built, CV can be used to locate and assess the quality of rebar lattices/mats [20, 21, 22, 23], specifically the spacing and linkage between rebar elements as a crucial parameter to their contribution to the structural integrity of bridges and buildings. While lattices have a characteristic regular structure, there also exists work on locating individual bars, e.g., in precast reinforced concrete [24]. During the lifetime of a building or other infrastructure, it is important to monitor structural health. Defects can either appear on the surface in the form of spallations and exposed rebars—as mentioned previously, or underneath the surface. Sub-surface structural health can be surveyed using CV techniques on Ground Penetrating Radar (GPR) data [25, 26, 27] as opposed to commonly used RGB cameras and 3D Light Detection and Ranging (LiDAR) sensors.
@@ 2.3 Construction robots
The previously mentioned related work focuses on inspection of structures and site monitoring and shows plenty of research activity in these areas. They provide a good basis for the adoption of autonomous systems, which rely on semantic understanding capabilities to navigate construction sites and carry out construction activities. Examples of such systems are excavators capable of autonomously building stone walls [28] or autonomous surface excavation [29]. Other examples are autonomous rebar tying robots [30, 31] and autonomous 3D concrete printing robots with rebar placement [32]. A pipeline detecting exposed rebars from RGB images and reconstructing a 3D model from stereo images for autonomous robotic placement of shotcrete is presented in [9, 33]. [34] developed a proof of concept showcasing the robotic application of shotcrete, and [35] shows an algorithm to identify and reconstruct steel rebar meshes. A thorough review of the state of construction robotics can be found in [36].
@@ 2.4 Datasets
Task-specific datasets are more tailored to solving a specific construction task. Table 1 gives an overview of datasets used in the works in Section 2.2 depicting rebars for various tasks. As the table shows, availability of data is low, which poses a barrier to the development of robust, deep learning-based solutions. Not every work makes a data availability statement, and more often than not, the data is only available upon request. Only [41] provides their data without any additional barriers. We want to overcome this barrier by releasing our dataset to the public, hoping that it will motivate other researchers to make also their data publicly available.
@@ 3.1 Overview
We have collected a dataset of 14,805 images captured mostly in construction environments. The main purpose of this dataset is to collect a broad set of images depicting exposed rebars in the context of shotcrete construction, aimed at autonomous inspection, repair, and construction of reinforced concrete structures. Our dataset contains scenes from various shotcrete construction tasks, such as the construction of ground support walls, tunnels, and culverts as well as the repair of piles and beams. This captures the diverse appearance of exposed rebars. The scenes consist of temporally ordered images, reflecting the robotic application this dataset targets.
@@
  • 1.
    Exposed rebars: These can be whole rebar lattices, i.e., in ground support wall construction, or partially exposed lattices and single bars in defective concrete.
  • 2.
    Persons: These will usually be construction workers in PPE.
  • 3.
    Cars: Those are regular-sized passenger cars.
  • 4.
    Trucks: This is usually heavy construction machinery or material delivery trucks.
Given the embedding of this work in the RobétArmé project  [^1] https://www.robetarme-project.eu/, which develops an autonomous robotic system for shotcrete inspection, repair and construction, we put our main focus on the exposed rebars class since it is important to identify these areas as they need to be filled with shotcrete. The other classes were chosen as sanity check classes, aimed at supervising the annotation process and empirically assessing the quality of the annotations.
In total, there are 54,115 instances, distributed as follows: 41,237 exposed rebars, 11,275 persons, 330 cars and 1273 trucks. Our dataset is a mix of self-collected data, for which the setup is described in Section 3.2, and YouTube videos.
@@ 3.2 Acquisition of data and labeling process
The scenes in the dataset stem from two sources: self-collected data we acquired during construction site visits and data from YouTube. The camera we used in our data collection visits is a Basler a2A1920-160ucBAS RGB camera fitted with a Basler C125-0418-5M-P f4mm lens. The camera’s exposure time is set to 5 ms, with automatic white balancing and the light source preset is set to Tungsten2800K. To avoid noisy images, we set the noise reduction and sharpness enhancement parameters to -10,000. We performed two data collection sessions for our self-captured data. The collection has been conducted in a hand-held manner, so as not to introduce any platform-specific characteristics into the data, possibly jeopardizing a general applicability of the dataset on different robotic platforms. A summary of the two recording sessions is included in the first two columns of Table 2.
Table 2: Metadata of the three constituent parts of ConRebSeg
Location Vester Sogade, 1601 Copenhagen, DK Langebro, 1219 Copenhagen, DK Web (YouTube)
@@
Short Description Balcony repair works with exposed rebars, both individual and in lattice format Repair of bridge piles and beams in challenging lighting conditions Mix of ground support walls, piles and beams, and tunnels and culverts
Environmental conditions outdoors, cloudy, naturally illuminated indoors, dark, artificially illuminated varying
Number of scenes 12 14 51
Number of frames 4911 7270 2624
Average frame rate 1.8 FPS 3.8 FPS N/A
Resolution (width/height) 1920x1200 1920x1200 varying
@@
Arranging construction site visits always relies on partners with connections to the construction industry, which can be a great barrier in the process, since not every site owner is willing to let outsiders in. That is why in addition to the self-collected data, we have also identified publicly accessible videos on YouTube depicting shotcrete construction activity, which has allowed us to rapidly grow the dataset. However, the frames of those videos are not directly included in the dataset, but rather their annotations together with a mapping to a specific frame of the video. This is due to legal uncertainties regarding the distribution of downloaded YouTube videos  [^2] In their service terms, YouTube states that ”You are not allowed to: […]1. access, reproduce, download, distribute, transmit, broadcast, display, sell, license, alter, modify or otherwise use any part of the Service or any Content except: […] (c) as permitted by applicable law;” and ”[…] access the Service using any automated means (such as robots, botnets, or scrapers) except: […] (c) as permitted by applicable law;[43]. We stipulate that the EU DSM directive [44] and its national implementation in Denmark [45] are fulfilling the conditions on which users are allowed to download data from YouTube as per their service terms.. Instead, we provide a way for users to obtain the data themselves in an automated manner for scientific/research purposes. Users should check the legality of this process in their respective jurisdictions. Once we have collected the data, we annotated them to obtain masks for the previously listed object classes. To make the annotation process privacy-preserving, we have treated all scenes with the face-blurring algorithm Deface [46], based on the CenterFace [47] face detector. Furthermore, we have treated all frames of the scenes gathered at Langebro (see Table 2) with histogram equalization and AI-denoising based on NAFNet [48] to alleviate the rather low-illumination, low-contrast frames. This facilitates the annotation process and avoids low-quality annotations due to invisible image features. However, we provide the original, but anonymized images in our dataset to give users the freedom to pre-process the frames according to their needs while maintaining the needed level of privacy. The experiments conducted in Section 4 use the raw version of the images without any level of anonymization, and we stipulate that the effect on the outcome of the experiments remains minimal.
@@
While classes such as person, truck, and car have some common sense associated with them, the ExposedBars class is rather domain-specific and not easy to describe. We set the following description for the annotation process: area with exposed rebars, where the tips of the bars should be polygon points and roughly outline the area. Along that, we gave some visual examples, as illustrated in Figure 1.
@@ 3.3 Statistics
Figure 2 also shows a clear dominance of the ExposedBars class, followed by the person class. We neglect this class imbalance as we are mainly interested in the ExposedBars category, given that there exists a plenitude of segmentation models trained on general-purpose datasets like MS COCO [49], Pascal VOC [50] or Cityscapes [51] for the remaining categories in our dataset. In total, there are 8570, 3543, and 2692 samples for the train, val, and test split respectively. Table 3 shows key statistics about the splits and their annotations.
@@ 3.4 Distribution and structure of the dataset
We provide a public repository that contains our data and other elements that facilitate viewing, exploring, and experimenting with the dataset. We use the FiftyOne dataset management framework [52] that provides end-users with a web interface for viewing samples and managing metadata such as train/test/val tags as well as other tags describing certain properties of samples. Our repository contains scripts that set up the dataset for the user, especially obtaining the data from YouTube, as we do not provide the frames directly for reasons specified in Section 3.2. In general, the data is organized as listed in Figure 4.
@@
The angle brackets in Figure 4 are placeholders and denote the meaning of the fields. dddd are decimal seconds, ns_epoch_timestamp denotes a timestamp in nanoseconds in Unix Epoch time (since Jan. 1, 1970), %04d means that frame numbers up to 1000 are zero-padded.
The YouTube frames are downloaded as whole videos and then are converted to frames, however, only every 200th frame is used since the informational value of the frames in between is minimal due to minimal variations of the depicted scene. For integrity purposes, we save a MD5 checksum for every sample in the dataset, which is saved as a metadata field in the dataset management tool. At import time, every frame is checked against this checksum to ensure the integrity of the dataset.